OpenStudy (anonymous):

Solve the equation algebraically 6x^2+7x=3

OpenStudy (anonymous):

6x^2+7x-3=0 x^2+7/6x-3/6=0 x^2+7/6x+49/144-49/144-1/2=0 (x+7/12)^2-(49/144+1/2)=0 (x+7/12)^2-121/144=0 (x+7/12)^2=121/144 (x+7/12)=+-(11/12) x=11/12-7/12=1/3 x=-11/12-7/12=-3/2

OpenStudy (anonymous):

How did you get + 49/144??

OpenStudy (anonymous):

to make it square you should add and minus b^2/4 to equation, that equation is in this form: x^2+bx+c=0 it's important that x^2 should not have any coefficient

OpenStudy (anonymous):

Got it, thanks!

OpenStudy (anonymous):

your welcome